• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name peter

    Log entry time 08:03:12 on August15,2003

    Entry number 110922

    keyword=Owl shift summary

    Shift Leader: Peter Monaghan
    Target Operator: Jaideep Singh
    3rd Person: Andrew Puckett

    00:00 Shift begins with elastic data being taken on kinematic
    point 9.03.

    00:15 MCC inform us that they are having a problem with one of the
    RF cavity; beam will be down for several minutes while they
    fix it.

    00:35 Beam returns.

    00:50 In rebooting and testing the left arm the run numbers got
    out of sync. Thus, run numbers 3541,3542 on the right arm
    do NOT exist.

    01:04 Set beam half-wave plate to be OUT.

    02:21 Finished elastic kinematics. Asked MCC for a controlled access
    to replace several fuses in transverse laser #3.

    02:24 Changing kinematics to 9.10; HRS-R/septum momentum = 1.076 GeV/c.

    03:10 Jaideep has finished working in the laser hut and we are
    waiting to go back to beam permit. The target half-wave plate
    is now OUT, the field direction is now 0 degrees, the
    target polarisation is 40.7%.

    03:15 Changing to the empty reference cell.

    03:43 Filled empty cell with nitrogen to 100 PSI.

    03:53 Changing to the polarised helium-3 target.

    04:05 After checking the T1 rates, we have increased the beam
    current to 10uA.

    04:15 Hall C bleedthrough measurement is 1.6% at 75uA delivered
    to hall C.

    04:20 Now inserting the beam half-wave plate.

    04:36 Rotating the target field direction to 270 degrees.

    05:05 Hall B are not ready to change the half-wave plate yet, so we
    will continue to take data with it IN. We are also going to
    reduce the beam current to 5uA to hopefully prolong the life
    and polarisation of the target.

    05:25 Now removing the beam half-wave plate.

    05:42 Finished kinematic point 9.10; now changing kinematics to
    9.11, with HRS-R/septum momentum = 0.995 GeV/c and septum
    current = 125.2A

    06:04 We have finished the kinematic change and are now starting
    9.11 with 3uA beam on the empty reference cell.

    06:22 Now fill the empty reference cell with nitrogen. We have
    discovered that the raster spot is rectangular and not square
    and so we are adjusting it appropriately - require a raster
    spot of 3.5 x 3.5 mm.

    06:44 Raster is now looking better. We are going to repeat the
    empty reference cell run for this 9.11 kinematic point. Note
    the cell has NOT yet been filled with nitrogen.

    07:00 Now filling the empty cell with nitrogen.

    07:05 NOTE: After analysing the latest empty reference cell runs
    we have seen that a large number of events pass all the cuts.
    The spectrum of events from the target appears very similar
    to the spectra given by the nitrogen filled reference cell.
    Further information can be found in the Halog entries by
    Jaideep.

    07:14 Now changing target to the polarised helium-3 target.

    07:19 MCC having problems with the machine and beam will be
    restored in a few minutes.

    07:44 MCC are continuing to have problems with a valve that is
    tripping out; they are still trying to fix the problem.

    07:48 New polarisation measurement: NMR = 40.4%; EPR = 39.4%

    07:57 Beam restored.

    -----------------------------------------------------

    Kinematic: 9.03 (Elastic)
    HRS-R/Septum Momentum = 1.142
    Septum Current = 143.5 A

    00:00 - 00:16 23539/3539 Pol. 3He - 180 hwp-IN 2.2M events 3uA beam

    00:36 - 00:49 3540 Pol. 3He - 180 hwp-IN 3M events 3uA beam
    Note: Left arm DAQ crashed and required rebooted; see note about
    concerning the run numbers.

    00:53 - 01:04 23543/3543 Pol. 3He - 180 hwp-IN 3M events 3uA beam

    Kinematic: 9.04 (Elastic)
    HRS-R/Septum Momentum = 1.142
    Septum Current = 143.5 A

    01:12 - 01:25 23544/3544 Pol. 3He - 180 hwp-OUT 3M events 3uA beam

    01:27 - 01:39 23545/3545 Pol. 3He - 180 hwp-OUT 3M events 3uA beam

    01:40 - 01:53 23546/3546 Pol. 3He - 180 hwp-OUT 3M events 3uA beam

    01.54 - 02:09 23547/3547 Pol. 3He - 180 hwp-OUT 3M events 3uA

    02:10 - 02:21 23548/3548 Pol. 3He - 180 hwp-OUT 3M events 3uA beam


    Kinematic: 9.10
    HRS-R/Septum Momentum = 1.076
    Septum Current = 135.5 A

    03:27 - 03:37 23549/3549 Empty Ref. Cell - hwp-OUT - 0 1uA beam 380k events.

    03:44 - 03:45 23550/3550 N2 ref. cell - hwp-OUT - 0 3uA current too high.

    03:47 - 03:53 23551/3551 N2 ref. cell - hwp-OUT - 0 1uA beam 1.3M events

    04:06 - 04:19 23552/3552 Pol. 3He - hwp-OUT - 0 10uA beam 2.6M events

    04:23 - 04:35 23553/3553 Pol. 3He - hwp-IN - 0 10uA beam 3M events

    04:52 - 05.05 23554/3554 Pol. 3He - hwp-IN - 270 10uA beam 3M events

    05:12 - 05:23 23555/3555 Pol. 3He - hwp-IN - 270 5uA beam 3M events

    05:28 - 05:41 23556/3556 Pol. 3He - hwp-OUT - 270 5uA beam 3M events


    Kinematic: 9.11
    HRS-R/Septum Momentum = 0.995
    Septum current = 125.2A

    06:07 - 06:13 23557/3557 Empty ref. cell - hwp-OUT - 0 3uA beam

    06:13 - 06:21 23558/3558 Empty ref. cell - hwp-OUT - 0 3uA beam 455k events

    06:30 - 06:42 3559 - 3561 runs are junk; used only to check the raster size.

    06:47 - 06:58 23562/3562 Empty ref. cell - hwp-OUT - 0 3uA beam 0.5M events

    07:04 - 07:14 23563/3563 N2 ref. cell - hwp-OUT - 0 3uA beam 2.2M